### v1

# Frozen Large Language Models Can Perceive Paralinguistic Aspects of Speech

Abstract
As speech becomes an increasingly common modality for interacting with large language models (LLMs), it is becoming desirable to develop systems where LLMs can take into account users’ emotions or speaking styles when providing their responses. In this work, we study the potential of an LLM to understand these aspects of speech without fine-tuning its weights. To do this, we utilize an end-to-end system with a speech encoder; the encoder is trained to produce token embeddings such that the LLM’s response to an expressive speech prompt is aligned with its response to a semantically matching text prompt where the speaker’s emotion has also been specified. We find that this training framework allows the encoder to generate tokens that capture both semantic and paralinguistic information in speech and effectively convey it to the LLM, even when the LLM remains completely frozen. We also explore training on additional emotion and style-related response alignment tasks, finding that they further increase the amount of paralinguistic information explicitly captured in the speech tokens. Experiments demonstrate that our system is able to produce higher quality and more empathetic responses to expressive speech prompts compared to several baselines.
Index Terms:
large language model, speech language model, speech emotion understanding, spoken dialogue interface
I Introduction
Large language models (LLMs) [1, 2, 3, 4] have become the foundation for many AI technologies today thanks to their strong natural language capabilities, as well as their ability to generalize to and perform tasks that they were not explicitly trained on. These characteristics, along with instruction tuning to align their behavior with human preferences [5, 6, 7], have led to systems that users can interact with seamlessly in a variety of settings.
Although the standard way of interacting with LLMs is via text-based prompting, speech is rapidly becoming a more common modality for this purpose. Speech contains paralinguistic cues that convey information such as emotion, emphasis, and focus [8], which makes it a richer form of communication than written text. Consequently, when developing speech-based interfaces to LLMs, it is becoming increasingly desirable to design systems where the LLM can take into account users’ emotions or speaking styles when providing responses to spoken prompts.
There has been a wide range of work on augmenting LLMs with speech and/or audio understanding capabilities, which can roughly be divided into two categories. The first category aims to adapt LLMs for solving speech and audio-related tasks, such as speech recognition, speech translation, or acoustic event recognition. These types of models may be specifically tailored to solve one or a small number of tasks [9, 10, 11], or they may be trained on a large number of tasks and datasets in an effort to learn general speech or audio understanding capabilities [12, 13, 14, 15, 16, 17]. However, they do not retain the original zero-shot instruction following capabilities of their base LLMs, and thus cannot be used for general purpose interactions with users. The second category aims to provide LLMs with the ability to understand and process speech inputs in the same way as text, essentially acting as a speech interface for prompting LLMs [18, 19, 20]. Most of these methods focus on conveying only the semantic information in speech to the LLM, disregarding paralinguistic information.
Recently, there have been several attempts to extend models in the second category to understand paralinguistics from speech. These works have approached the task by concatenating additional emotion or paralinguistic features to the input prompt tokens [21, 22] or by training a speech encoder to produce token embeddings that convey emotional information to the LLM [23]. However, due to the relative scarcity of emotion-labeled speech data, these models are often trained using synthetic data created using expressive text-to-speech synthesis, which may limit their ability to generalize to natural human speech. Furthermore, they all use low-rank adaptation (LoRA) [24] to fine-tune their LLMs so that they can process the transformed speech or emotion tokens. While LoRA enables efficient adaptation of LLMs for specific use cases, it can make a model difficult to use in settings that it was not fine-tuned for, and it in general does not guarantee preservation of the full capabilities of the base model. Having a modality adapter for speech that can simply be attached on top of an LLM without changing any of its parameters is a much more straightforward approach for many real-world deployment scenarios.
In this work, we study the capabilities of an LLM to understand emotions and paralinguistic information in speech signals without any kind of fine-tuning. We utilize an end-to-end system that combines a frozen text LLM (Llama 3 8B Instruct [4]) with a speech encoder that converts spoken inputs into token embeddings that the LLM can process. The encoder is trained to produce speech tokens that convey both semantic and paralinguistic information to the LLM, which the LLM can take into account when responding to a spoken prompt. To do this, we use a speech emotion recognition (SER) dataset with transcriptions; our training framework guides the speech encoder to generate tokens such that the LLM’s response to an expressive speech prompt is aligned with its response to a semantically matching text prompt where the speaker’s emotion or speaking style has been specified. We refer to this process as alignment against emotion-conditioned responses. We also explore training the system on other emotion-related response alignment tasks, specifically speech captioning and SER, and find that this further improves the encoder’s ability to capture paralinguistic information in the speech tokens.
Interestingly, despite only being trained on causal language modeling loss, our framework produces speech tokens that can effectively convey the emotion or speaking style of a spoken input to the LLM. The resulting system, which we call SpeechEmotionLlama, is able to generate higher quality and more empathetic responses to expressive speech prompts compared to versions trained without emotion-conditioned response alignment, as well as a cascade of a standalone SER model and the non-emotion-conditioned baseline. Our findings open up interesting potential research directions on utilizing frozen LLMs to perform other speech or audio-related tasks.
II Preliminaries
II-A Behavioral alignment between speech and text
Our work builds on prior research on augmenting frozen LLMs with the ability to process speech inputs [18]. The overall model is an end-to-end system that consists of two components: an LLM and a speech encoder that converts audio into speech tokens that the LLM can process. The key idea here is behavioral alignment between speech and text inputs to an LLM; that is, for two prompts that convey the same information, the LLM should provide the same response regardless of whether the prompt is in speech or text. Training such a system requires paired speech-text samples of spoken utterances and their transcripts. Specifically, the speech encoder is trained to produce token embeddings such that the LLM’s response to a spoken prompt matches its response to the paired text prompt, which is treated as the ground truth. This alignment may be done in multiple ways, such as language modeling against the ground truth response [18, 19] or knowledge distillation from the text input [20, 25].
Here, we use causal language modeling to align the two modalities, conditioning the prediction of the response text tokens on the input speech tokens. Formally, let the full token embedding sequence fed into the model during training be $\mathbf{x}=(x_{0},...,x_{t},x_{t+1},...x_{T})$, where the first $t$ elements correspond to the prompt, including the speech tokens and any tokens used in the LLM’s chat structure: $(x_{0},...,x_{t})$. The remaining tokens correspond to the LLM’s ground truth response to the paired text prompt: $(x_{t+1},...,x_{T})$. Then, the training loss is:
$\mathcal{L}_{\mathrm{LM}}=-\sum_{i=t+1}^{T}\log p(x_{i}|x_{<i}).$ (1)
Note that the loss is computed only for token indices corresponding to the response portion of the sequence, $[t+1,T]$.
II-B System architecture
We use Llama 3 8B Instruct [4] as our LLM, which remains entirely frozen throughout all of our experiments. For the speech encoder, we use a 24-layer Conformer-based [26] model with 1B parameters. It takes as input 80-dimensional mel-spectrograms which are computed from 16kHz audio using a window length of 25ms and a shift of 10ms. Every four spectrogram frames are stacked to temporally compress the input sequence by a factor of 4. The stacked output of the Conformer layers is fed into a 100M parameter adapter module composed of a convolution layer, a rotary Transformer layer, and a linear layer which maps the output dimension to that of the LLM’s input embeddings (4096-dim). The convolution layer in the adapter further compresses the features by a factor of 2 to produce speech tokens every 80ms.
The Conformer layers of the speech encoder are first pre-trained using the BEST-RQ method [27] on around 15M hours of multilingual speech data. Then, they are jointly optimized along with the (randomly initialized) adapter module on the behavioral alignment criterion described in Section II-A. This is done using a mixture of 230K hours of automatic speech recognition (ASR), 90K hours of automatic speech translation (AST), and 25K hours of spoken dialogue data. All datasets described above are internal.
After these initial training stages, the encoder is able to produce speech token embeddings that can be processed by the LLM as if they were no different from standard text token embeddings, essentially acting as a speech interface to leverage all of the LLM’s capabilities. The speech tokens can either be used on their own or interleaved with text tokens as part of a multimodal prompt. More specific details on the speech encoder’s architecture, pre-training, and alignment process can be found in [4].
Figure 1: Behavioral alignment against emotion-conditioned responses. The speech encoder is trained to produce token embeddings such that the LLM’s response to an expressive speech prompt matches its response to a semantically paired text prompt where the speaker’s emotion or speaking style has also been specified.
III Proposed Method
III-A Behavioral alignment against emotion-conditioned responses
The framework described above focuses on aligning the responses of the LLM according to the semantic information in speech and text inputs. Therefore, it does not enable the speech encoder to capture or convey paralinguistic cues in a spoken utterance; speech inputs that convey the same words would be expected to elicit very similar responses, regardless of the emotion or style in which they were spoken. In the rest of this work, we aim to address the question: Is it possible to use the behavioral alignment paradigm to enable a frozen LLM to understand paralinguistic cues from speech inputs?
We propose an approach that performs LLM response alignment between the text and speech modalities similarly to above, but where the ground truth text response has been generated with the addition of an expressive emotion or style tag. This leverages a key observation about Llama 3’s behavior, in which the LLM responds differently to the same text input if an emotion or speaking style is also specified as part of the prompt. This behavior can be made even more explicit by adding a system prompt instructing the LLM to respond as if it is being spoken to in that way. We refer to this process as behavioral alignment against emotion-conditioned responses. Fig. 1 shows examples of emotion-conditioned ground truth responses and how they fit into the behavioral alignment training paradigm.
Specifically, we now utilize a speech emotion recognition (SER) dataset of paired speech-text samples that also contain expressive speaking style labels for the speech utterances. We generate ground-truth responses to text inputs using the following prompt template:
"Respond as if you are a natural conversation companion.
Do not apologize, and do not refer to the fact that you are a large language model or an AI. If the user input
indicates an emotion, respond as if you are being spoken to in that emotion: <emotion> {prompt_transcript}"
Then, starting from the model weights resulting from the training procedure in Section II-B, we further fine-tune the speech encoder on response alignment against these emotion-conditioned responses. Here, we do not prepend any kind of system prompt before the speech tokens, unlike how the ground-truth text response is generated. This is done in order to more fundamentally alter the behavior of the encoder and make it capture paralinguistic information in the speech tokens, rather than create any dependencies on a specific system prompt. We refer to the resulting end-to-end system as SpeechEmotionLlama.
Our method shares some similarities with the emotional response continuation task in BLSP-Emo [19] and the speech emotion captioning framework in SECap [28]. However, unlike BLSP-Emo, we do not use any kind of system prompt before the speech tokens asking the LLM to take into account the emotional tone, and we do not use LoRA to fine-tune its weights. Meanwhile, SECap focuses specifically on speech emotion captioning, while we aim for a general speech interface that enables users to interact with LLMs for any purpose.
III-B Additional tasks
In addition to the emotion-conditioned response matching task described above, we also train on several other types of response alignment tasks, which aim to improve the ability of the speech encoder to capture semantic and/or paralinguistic information in its tokens. This is achieved by prepending specific text prompts to the speech token embeddings before feeding them into the LLM.
ASR. We ask the LLM to perform ASR using the following prompt:
"Repeat after me in English: {input}"
The speech tokens are used as “{input}”, and the target text is the transcript of the utterance. This task reflects the ASR training method described in [4] and is meant to help the speech encoder better capture the semantic information in spoken inputs.
Speech captioning. We ask the LLM to perform speech captioning using the following prompt:
"You will hear a speech utterance spoken by a person. In natural language, provide a caption that faithfully describes the scenario in which the speech utterance is being spoken. Pay attention to the spoken content, and make sure to incorporate any emotions expressed in the speech into the caption as needed: {input}"
Similarly to response matching in Section III-A, we generate ground truth captions by using “<emotion> {prompt_transcript}” as “{input}”. During training, we use the speech tokens as “{input}” and the ground truth caption as the target text.
SER. We ask the LLM to perform SER using the following prompt:
"Identify the emotional tone of the utterance that is provided. Select from the following options: neutral, angry, disgusted, fearful, happy, sad, surprised, bored, confused, interested, sarcastic, laughing, projected, sleepy, whisper: {input}"
As with the other tasks, the speech tokens are used as “{input}”. The target text is the text of the emotion/style tag for the given utterance, which is one of the 15 classes shown in the text prompt (see Section IV-A for more details on the data).
IV Experimental Setup
IV-A Datasets
We use a combination of 3 internal English speech datasets with emotion/style labels. Two of the datasets consist of expressive speech from voice actors who are native English speakers; they comprise 110 hours of speech from 10 speakers and 82 hours of speech from 10 speakers, respectively. Of these, the first is a superset of the Expresso dataset [29]. The third dataset is an expressive speech corpus of around 2300 hours spoken by 21K speakers; the speakers are not voice actors, and are made up of a mixture of native and non-native English speakers. All three datasets were filtered to remove data from geographically sensitive areas. The data encompasses 15 style classes: neutral, angry, disgusted, afraid, happy, sad, surprised, bored, confused, interested, sarcastic, laughing, projected, sleepy, and whisper.
We use most of the data for training, but hold out 1 speaker each from the first two datasets and 5% of the third dataset for validation (about 134 hours total). We also create a test set of 2157 samples spoken by 1 additional held out speaker from each of the first two voice actor datasets (2 speakers). The test set was constructed in this way because a large portion of the first two datasets consists of speech utterances where the same content is spoken in multiple emotions or styles, or where the content does not necessarily reflect the emotion being conveyed. Evaluating on these samples allowed us to better evaluate models’ abilities to process paralinguistic information in speech while controlling for semantic content.
TABLE I: Performance of various model configurations in terms of metrics measuring response quality to expressive speech prompts and on speech emotion recognition (SER). We also include 95% confidence intervals for the two LLM-annotated scores.
Speech Encoder Fine-Tuning Tasks ROUGE-1,2,L BERTScore LLM Score SER
Response ASR Caption SER Quality Emotion
Baseline 1 (not fine-tuned) 0.314, 0.112, 0.213 94.64 5.46 ± 0.13 4.13 ± 0.13 20.45
Cascade: SER + Baseline 1 0.325, 0.122, 0.224 94.64 5.58 ± 0.13 4.58 ± 0.14 80.81
Baseline 2 (fine-tuned) 0.375, 0.145, 0.245 94.73 7.57 ± 0.07 5.59 ± 0.11 13.89
Cascade: SER + Baseline 2 0.381, 0.151, 0.250 94.78 7.63 ± 0.07 5.83 ± 0.11 80.81
SpeechEmotionLlama v1 0.437, 0.214, 0.309 94.76 7.79 ± 0.07 7.56 ± 0.10 41.26
SpeechEmotionLlama v2 0.439, 0.218, 0.312 94.75 7.79 ± 0.06 7.45 ± 0.10 59.95
SpeechEmotionLlama v3 0.446, 0.224, 0.318 94.79 7.74 ± 0.07 7.54 ± 0.10 81.51
IV-B Baselines
We compare against several different system configurations: one where the speech encoder is not fine-tuned after the training described in Section II-B, one where it is fine-tuned on the new datasets using ASR and response alignment against non-emotion-conditioned responses, and versions where the speech encoder is fine-tuned on the new datasets using different subsets of the four alignment tasks.
We also experiment with incorporating emotion/style information into the first two baselines by cascading them with a standalone SER model. The SER model is a time and layer-wise transformer (TLTR) [30] trained on top of the first baseline’s frozen speech encoder; it uses temporal and layer-wise attention to aggregate the encoder’s intermediate layer representations and perform classification. The model achieves 80.81% unweighted accuracy on the 15-way classification task on our test set. For the cascaded systems, we predict the emotion/style class of the input speech, convert it to a string with delimiters (e.g., “<happy>”), tokenize the string, and prepend the resulting token embeddings to the speech token embeddings before feeding them into the LLM, matching the format when the ground truth emotion-conditioned responses are generated (see Section III-A).
IV-C Training and evaluation
Training was done on 64 NVIDIA A100 80GB GPUs with a batch size of 4 per GPU. We used the Adam optimizer [31] with $\beta_{1}=0.9$, $\beta_{2}=0.98$ and a learning rate of 3e-5. All model variants were trained for 5 epochs with 20,000 warm-up steps, and were optimized for an equal number of steps on each of the tasks they were trained on.
We evaluate models by measuring the quality of their responses to expressive speech prompts. We use ROUGE [32] and BERTScore [33] to compare the system outputs against the ground truth emotion-conditioned responses. We also employ Llama 3 Instruct 70B [4] as an evaluator to score the responses generated by the different systems on a scale from 1-10 in terms of 1) overall quality, helpfulness, and relevance, and 2) how well the response took into account the user’s emotion and/or how empathetically the model responded.
TABLE II: Responses of Baseline 2 and SpeechEmotionLlama v3 to the speech prompt “I can’t believe I ripped my pants when I came over the wall though” spoken in the “laughing” style. Note that the baseline does not perceive or acknowledge the laughing nature of the speech in its response, while SpeechEmotionLlama does.
Model Response
Baseline 2 “Oh no, that’s frustrating! It’s always a bummer when something like that happens. Did you get a good laugh out of it or was it more of a "oh no, not again!" moment? And what was going on when you came over the wall? Were you trying to jump or just trying to get to the other side?”
v3 “Oh no, that’s hilarious! I can just imagine the look on your face! Did you have to do a quick fix or did you just laugh it off and keep going? What was going on that made you jump over the wall so suddenly? Was it a dare or a sudden impulse?”
V Results
V-A Responses to expressive speech prompts
Table I shows the performance of the various models in terms of the quality of their responses to expressive speech prompts. Baseline 2 obtained much better scores than Baseline 1, thanks to having been fine-tuned on in-domain data. However, its emotion understanding score was still quite low; this is unsurprising, since its training should not have allowed it to learn or utilize any paralinguistic information from speech inputs. Cascading both baselines with the SER model resulted in marginal gains in terms of response quality and emotion understanding, but the improvements were quite small.
Modifying training to perform alignment against emotion-conditioned responses resulted in large improvements in terms of all metrics. ROUGE, BERTScore, and LLM-annotated quality scores all increased for SpeechEmotionLlama v1, but perhaps the most notable improvements were in emotion understanding, which improved by over 1.7 points compared to the cascaded SER + Baseline 2 system. Adding additional style and emotion-related response alignment tasks (v2 and v3) did not lead to significant changes to either of the LLM-annotated scores, but provided slight improvements to ROUGE and BERTScore. These results indicate that our proposed training framework enabled SpeechEmotionLlama to take into account the emotion or speaking style of a speech prompt and respond accordingly, unlike baselines that were not trained on emotion-conditioned responses. Table II shows examples of responses from Baseline 2 and SpeechEmotionLlama v3 to an expressive speech prompt demonstrating this.
V-B Speech emotion recognition
To more explicitly examine how well each system’s speech tokens convey emotion/style information to the LLM, we prompted the different model variants to perform SER by prepending the same prompt as used in the SER training task before the speech tokens. Then, we checked whether the response generated by the LLM matched the emotion/style label. The rightmost column of Table I shows these results; the only exceptions are the cascade models, for which we list the performance of the standalone SER model. As expected, the two baselines performed quite poorly. Fine-tuning on emotion-conditioned response alignment resulted in a significant increase in performance, and adding speech captioning and SER each resulted in further improvements. Surprisingly, SpeechEmotionLlama v3 even outperformed the standalone SER model, achieving 81.51% accuracy. This further demonstrates that our training framework allows the speech encoder to effectively convey paralinguistic information to the LLM in a way that allows the LLM to make use of that information.
Figure 2: t-SNE plots of temporally pooled speech token embeddings from (a) Baseline 2 and (b) SpeechEmotionLlama v3. Colors denote the emotion/style tag of the utterances.
V-C Analysis of speech tokens
We performed extra analyses to study the amount of paralinguistic information encoded in the different systems’ speech tokens. We first trained linear probes on temporally pooled speech tokens from each model variant to perform the 15-way SER task. The two baselines obtained 25.08% and 31.20% accuracy, while SpeechEmotionLlama v1, v2, and v3 obtained higher accuracies of 68.15%, 73.15%, and 78.40%, respectively. We also performed a t-SNE [34] analysis of the mean-pooled speech tokens. Fig. 2 displays the resulting plot, which shows that the speech tokens from SpeechEmotionLlama are clustered much more clearly by emotion/style tag compared to the baseline. These results demonstrate that training on more emotion-related response alignment tasks produces token embeddings that encode more discriminative paralinguistic information from speech.
VI Conclusion
We study the potential of an LLM to understand paralinguistic aspects of speech without any fine-tuning of its weights. Building on prior work that equips frozen LLMs with the ability to process speech inputs, we propose a novel training paradigm called behavioral alignment against emotion-conditioned responses. Our framework enables a speech encoder to generate token embeddings that capture both semantic and paralinguistic information from a spoken utterance and effectively convey it to the LLM, even when the LLM remains completely frozen. The LLM, in turn, is able to effectively utilize the paralinguistic information in the speech tokens to provide higher quality and more empathetic responses to expressive speech prompts. Our work opens up new possibilities for developing interfaces to LLMs that can leverage their capabilities for other kinds of speech or audio tasks without the need for any fine-tuning.